owl walk through, 9Dec99

Marj Corcoran (corcoran@physics.rice.edu)
Thu, 9 Dec 1999 05:06:50 -0600 (CST)

Shift Summary for owl shift, 9/Dec/99

On Shift: Marj (walked through twice)

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: accelerator still off

========================================================================
Runs ended this shift:

Run 151655 ongoing, cosmics

Tapes brought to FCC this shift: None

========================================================================
Accesses during shift:

open access

========================================================================
Overview of shift:

On 2AM walkthrough, cosmic run is ongoing. Val is complaining about
gain drift in CsI.

VolMon-CsI button on SDAQ screen is red, due I think to ETOT being off as
per sticky note on the sdaq screen.

Note that DC3 pressure reading upstairs is low. It reads about .003 inches
of water, but all chambers should be at about 0.04 or 0.05 inches of
water. The reading downstairs on the magnehelic is OK (0.05), so the
pressure gauge upstairs is probably out of calibration. Low reading
on DC3 began in October.

5AM walkthrough, all still OK. Gas checklist done. Cosmics still ongoing.
Val still complaining about gain drift.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

CsI gain drift
--------------------------------------------------------